OpenStudy (anonymous):

A comfortable summer day is 77°F. What is this on the Celsius scale?

OpenStudy (anonymous):

F = (9/5)C + 32.

OpenStudy (anonymous):

F = (9/5)C + 32. =(9/5)C+32 77 = (9/5)C +32 simplifying we find C=25 celcius